OpenStudy (anonymous):

HELP! functions? The velocity "v" if the airstream through an average-sized throat is related to the radius "r" of the throat (in cm) by the function: v(r)=3.2(1-r)r^2, 0.5<=r<=1 Determine value r for which v is a maximum? Please help me :) <= means less than or equal to

OpenStudy (cwrw238):

use differentiation v(r) = 3.2r^2 - 3.2r^3 v' =6.4r - 9.6r^2 for maxm r : 6.4r - 9.6r^2 = 0 3.2r(2 - 3r) = 0 3r = 2 r = 2/3 or 0.67 cm to 2 dec places

OpenStudy (cwrw238):

* for maxm v:

OpenStudy (cwrw238):

ok ? are you familiar with calculus?

OpenStudy (anonymous):

no, i'm only in algebra 2... is there any way to do this with material i would know?

OpenStudy (anonymous):

someone help please! :)